How do I find the slope of the line containing the points (-2,1) and (4,-3)?

Answer by mananth(16946)

x1 y1 x2 y2
-2 1 4 -3

m= (y2-y1)/(x2-x1)
m=(-3-1)/(4-(-2))
=-4/6
=-2/3